# Craftax: A Lightning-Fast Benchmark for Open-Ended Reinforcement Learning

### v1

@@ 2.4 Open-Ended Learning
Unsupervised Environment Design Unsupervised environment design (UED) is a paradigm in RL where an adversary proposes environments configurations (referred to as levels) for an agent to train on (Dennis et al., 2020; Jiang et al., 2021a; Parker-Holder et al., 2022). One popular approach is to let the adversary choose levels that maximise the agent’s regret—corresponding to the difference in performance between the optimal policy and the agent. Minimax regret tends to incentivise the adversary to generate a curriculum of initially easy, then increasingly challenging environments. Training on these environments empirically improves the generalisation of agents.
@@ 3.2 Craftax: An Extension of Crafter with NetHack-Like Mechanics in JAX
Multiple Floors While in Crafter the player is confined to a single 64x64 grid, Craftax contains 9 unique procedurally generated floors, including caves, dungeons, fire and ice floors and a final boss floor. The player can descend and ascend through the world by finding the ladders that connect adjacent floors. Each floor contains distinct challenges in the forms of different terrain generation, enemies and required skills, necessitating deep exploration and generalisation. While each floor is unique, the basic game mechanics carry across them, so an agent can partially transfer knowledge as it ventures deeper into the world.
@@
Combat We overhaul the combat system with damage and defence split into physical, fire and ice categories. As well as being able to craft higher-tier weapons, the player can use a bow (found in a random chest on the first dungeon floor) for ranged attacks, and read books (also found in dungeon chests) to learn offensive spells to cast. The player can also craft armour to protect themselves from damage (essential when descending to the lower floors). As well as increasing the general complexity, the diversity in combat furthers the in-context learning element provided by the procedural level generation – an agent that stumbles upon a strong weapon or armour piece should suitably change its strategy.
New Creatures The 3 creatures present in Crafter (cow, zombie and skeleton) have been increased to 19 different creatures, each with unique effects, behaviours and techniques required to defeat them. For instance, the Fire Elementals present on floor 7 either require an appropriate enchantment on the player’s sword or magic to defeat, while beating the aquatic enemies on floor 6 typically requires ranged weaponry.
@@
Attributes The player is rewarded with an experience point every time they descend to a new floor. These points can be spent to increase the players dexterity, strength or intelligence. Dexterity makes the player more capable with a bow, as well as allowing them to go longer between eating, drinking and sleeping. Strength increases the players health and melee damage, while intelligence increases the players magical powers. The assignment of experience points can drastically alter the players abilities and have a large influence on the strategy an agent should use, testing the agents long term reasoning capabilities.
@@ 3.3 RL Environment Interface
Figure 3: Rewards on Craftax-1B for PPO, PPO-RNN, ICM and E3B. Each algorithm is run for 1 billion timesteps with 10 seeds. The shaded area denotes 1 standard error.
@@ 4.1 Exploration Baselines
Intrinsic Curiosity Module The first baseline is a popular exploration method, the Intrinsic Curiosity Module (ICM) (Sukhbaatar et al., 2018). ICM learns a world model and provides intrinsic reward proportional to the error of this world model, incentivising the agent to visit areas of the environment which are badly modelled and therefore probably less visited. The latent space of this world model is constructed in such a way as to ignore aleatoric uncertainty in the environment, in order to focus exploration and avoid the so called ‘Noisy TV problem’ (Burda et al., 2018).
Exploration via Elliptical Episodic Bonuses Our second exploration baseline is the E3B algorithm (Henaff et al., 2022). Rather than considering exploration over the entire training process, E3B formulates its intrinsic reward to encourage exploration within each episode. An ellipse is fitted around the agents current trajectory in latent space and it is rewarded for moving far away from this ellipse. This approach has been shown to outperform global exploration methods, especially in non-singleton (i.e. procedurally generated) environments (Henaff et al., 2022, 2023).
@@ 4.2 UED Baselines
We benchmark variations of two common UED algorithms, PLR and ACCEL andalso compare to Domain Randomisation (Jakobi, 1997; Tobin et al., 2017, DR) , which corresponds to training on randomly generated levels as usual.
@@
In this section, we provide results for the two proposed benchmarks using the previously discussed algorithms. Hyperparameter tuning is detailed in Appendix E.
@@ 4.3 Craftax-1B
The returns for the evaluated algorithms on Craftax-1B are summarised in Figure 3 with the final achievement yields split by difficulty shown in Figure 4. We report reward as a percentage of the maximum achievable reward (obtained by completing every achievement) of 226 for readability. Fine-grained achievement results are shown in Appendix F.1 with a selected set of achievements highlighted in Figure 5. We see that PPO and PPO-RNN both quickly master most of the basic achievements, with memory helping to push progress further into the intermediate category. Interestingly, neither of the two tested exploration methods improved performance and E3B in fact significantly reduced the reward. We hypothesise this is because the reward structure is dense enough for RL methods without intrinsic reward to learn on, with the intrinsic reward potentially serving to distract the agent. None of the agents make any progress in the harder achievement categories.
@@ 4.4 Craftax-1M
Figure 6: Rewards on Craftax-1M for PPO, PPO-RNN, ICM and E3B. Each algorithm is run for 1 million timesteps with 10 seeds. The shaded area denotes 1 standard error.
@@ 6 Conclusion
We present Craftax, a blazing fast environment filled with complex mechanics that cannot be solved by existing RL algorithms. We hope that Craftax will facilitate research into areas including exploration, continual learning, generalisation, skill acquisition and long term reasoning. We believe an agent that could solve Craftax would represent a substantial step forward for the field and look forward to what the community develops using the benchmark.
@@ D.3 World Generation
Figure 11: Craftax: Floor 1 - Dungeons
Figure 12: Craftax: Floor 2 - Gnomish Mines
Figure 13: Craftax: Floor 3 - Sewers
@@
Figure 15: Craftax: Floor 5 - Troll Mines
Figure 16: Craftax: Floor 6 - Fire Realm
Figure 17: Craftax: Floor 7 - Ice Realm
@@ E.1 Craftax-1B
The hyperparameters and considered values for PPO on Craftax-1B are shown in Table 7. Each run was performed for 1 billion timesteps with 1 seed. We started from a set of values picked informally and then tuned each hyperparameter individually. These hyperparameters were then taken as the baseline for the other methods, which all build off of the PPO implementation. For PPO-RNN, we kept the same hyperparameters as PPO, with the hidden size being equal to the MLP layer size. The hyperparameters for ICM and E3B are shown in Tables 8 and 9 respectively.
@@ E.2 Craftax-1M
The hyperparameters for Craftax-1M were tuned more thoroughly, as the experiments were very fast to run. We performed a random search over the PPO hyperparameters to obtain a baseline PPO implementation, before then running a random search for ICM and E3B. The hyperparameters for PPO, ICM and E3B are shown in Tables 10, 11 and 12 respectively.
@@ G.2 Additional UED Results
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption

### v2

@@ 2.4 Open-Ended Learning
Unsupervised Environment Design Unsupervised environment design (UED) is a paradigm in RL where an adversary proposes environments configurations (referred to as levels) for an agent to train on (Dennis et al., 2020; Jiang et al., 2021a; Parker-Holder et al., 2022). The adversary is rewarded for choosing levels that maximise the agent’s regret, defined as the difference in return between the current and optimal agent. This has been empirically shown to automatically induce a curriculum of progressively harder levels that aid the performance and generalisation properties of the learned agent. Different UED algorithms require different levels of access to the underlying environment state, ranging from simply being able to repeat seeds (Jiang et al., 2021a) to directly editing the levels (Parker-Holder et al., 2022). Due to the functional nature of Craftax necessitated by JAX, the entire environment state is exposed as a single object, making UED methods easy to apply.
@@ 3.2 Craftax: An Extension of Crafter with NetHack-Like Mechanics in JAX
Multiple Floors While in Crafter the player is confined to a single 64x64 grid, Craftax contains 9 unique procedurally generated floors, including caves, dungeons, fire and ice floors and a final boss floor. The player can descend and ascend through the world by finding the ladders that connect adjacent floors. Each floor contains distinct challenges in the forms of different terrain generation, enemies and required skills, necessitating deep exploration and generalisation. While each floor is unique, many game mechanics are shared between them and, on a meta-level, exploration strategies that worked on earlier floors (for instance moving adjacent to a block and trying different actions to figure out its characteristics) will also work on later floors. In this way we hope to not only facilitate generalisation across different procedurally generated worlds but also generalisation of the exploration strategy through time over the learning process.
@@
Combat We overhaul the combat system with damage and defence split into physical, fire and ice categories. As well as being able to craft higher-tier weapons, the player can use a bow (found in a random chest on the first dungeon floor) for ranged attacks, and read books (also found in dungeon chests) to learn offensive spells to cast. The player can also craft armour to protect themselves from damage (essential when descending to the lower floors). As well as increasing the general complexity of the environment, the diversity in combat furthers the in-context learning element provided by the procedural level generation –- an agent that stumbles upon a strong weapon or armour piece should suitably change its strategy. This further extends the exploration problem as, by design, there should not be one fixed strategy (for instance, always putting experience points into strength and defeating enemies with melee attacks) that works on every level, meaning that an agent will have to explore a diverse range of strategies to achieve consistently high return.
New Creatures The 3 creatures present in Crafter (cow, zombie and skeleton) have been increased to 19 different creatures, each with unique effects, behaviours and techniques required to defeat them. For instance, the Fire Elementals present on floor 7 either require an appropriate enchantment on the player’s sword or magic to defeat, while beating the aquatic enemies on floor 6 typically requires ranged weaponry. This diversity of creatures further enhances the exploration problem, as no two are the same and strategies for dealing with one may entirely fail for others.
@@
Attributes The player is rewarded with an experience point every time they descend to a new floor. These points can be spent to increase the players dexterity, strength or intelligence. Dexterity makes the player more capable with a bow, as well as allowing them to go longer between eating, drinking and sleeping. Strength increases the players health and melee damage, while intelligence increases the players magical powers. The assignment of experience points can drastically alter the players abilities and have a large influence on the strategy an agent should use, testing the agents long term reasoning capabilities, as successful runs typically take tens of thousands of frames.
@@ 3.3 RL Environment Interface
Figure 3: Rewards on Craftax-1B for PPO, PPO-RNN, ICM, E3B and RND. Each algorithm is run for 1 billion timesteps with 10 seeds. The shaded area denotes 1 standard error.
@@ 4.1 Exploration Baselines
Random Network Distillation We use the popular exploration baseline Random Network Distillation (RND) (Burda et al., 2018) as our first method for global intrinsic reward. RND works by training a network that is given observations to match a network of the same architecture that was randomly initialised. Intuitively, the error of this network will be high on unseen observations, as the network will not yet have been trained to match the random output. This error is then used as an additional intrinsic reward that incentivises the agent to explore states with novel observations.
Intrinsic Curiosity Module Our second baseline is the Intrinsic Curiosity Module (ICM) (Sukhbaatar et al., 2018). ICM learns a world model and provides intrinsic reward proportional to the error of this world model, incentivising the agent to visit areas of the environment which are badly modelled and therefore probably less visited. The latent space of this world model is constructed in such a way as to ignore aleatoric uncertainty in the environment, in order to focus exploration and avoid the so called ‘Noisy TV problem’ (Burda et al., 2018).
Exploration via Elliptical Episodic Bonuses Our third and final exploration baseline is the E3B algorithm (Henaff et al., 2022). Rather than considering exploration over the entire training process, E3B formulates its intrinsic reward to encourage exploration within each episode. An ellipse is fitted around the agents current trajectory in latent space and it is rewarded for moving far away from this ellipse. This approach has been shown to outperform global exploration methods, especially in non-singleton (i.e. procedurally generated) environments (Henaff et al., 2022, 2023).
@@ 4.2 UED Baselines
We assess PLR and ACCEL and compare them to Domain Randomisation (Jakobi, 1997; Tobin et al., 2017, DR), which corresponds to uniformly sampling randomly generated levels, as per usual. We use JaxUED (Coward et al., 2024) as the base implementation for all our UED experiments.
@@
In this section, we provide results for the two proposed benchmarks using the previously discussed algorithms. We use JaxUED (Coward et al., 2024) for all UED experiments, and hyperparameter tuning is detailed in Appendix E.
@@ 4.3 Craftax-1B
The returns for the evaluated algorithms on Craftax-1B are summarised in Figure 3 with the final achievement yields split by difficulty shown in Figure 4. We report reward as a percentage of the maximum achievable reward (obtained by completing every achievement) of 226 for readability. Fine-grained achievement results are shown in Appendix F.1 with a selected set of achievements highlighted in Figure 5. We see that PPO and PPO-RNN both quickly master most of the basic achievements, with memory helping to push progress further into the intermediate category. Interestingly, none of the tested exploration methods improved performance and E3B in fact significantly reduced the reward. We hypothesise this is because the reward structure is dense enough for RL methods without intrinsic reward to learn on, with the intrinsic reward potentially serving to distract the agent. None of the agents make any progress in the harder achievement categories.
@@ 4.4 Craftax-1M
Figure 6: Rewards on Craftax-1M for PPO, PPO-RNN, ICM, E3B and RND. Each algorithm is run for 1 million timesteps with 10 seeds. The shaded area denotes 1 standard error.
@@ 6 Conclusion
We present Craftax, a blazing fast environment filled with complex mechanics that cannot be solved by existing RL algorithms. We hope that Craftax will facilitate research into areas including exploration, continual learning, generalisation, skill acquisition and long term reasoning. We believe an agent that could solve Craftax would represent a substantial step forward for the field and look forward to seeing what the community uses the benchmark to develop.
Acknowledgements
We would like to thank Alex Goldie and Sebastian Towers for their invaluable contributions to game design, our anonymous reviewers for their useful feedback and to the open-source software community who have already contributed numerous improvements to the codebase.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ E.1 Craftax-1B
The hyperparameters and considered values for PPO on Craftax-1B are shown in Table 7. Each run was performed for 1 billion timesteps with 1 seed. We started from a set of values picked informally and then tuned each hyperparameter individually. These hyperparameters were then taken as the baseline for the other methods, which all build off of the PPO implementation. For PPO-RNN, we kept the same hyperparameters as PPO, with the hidden size being equal to the MLP layer size. The hyperparameters for ICM, E3B and RND are shown in Tables 8, 9 and 10 respectively.
@@
Hyperparameter Considered Values Value
Intrinsic Reward Coefficient $\beta$$\{0.01,0.1,1.0\}$ $1.0$
RND Output Dimension $\{4,32,256,2048\}$ $256$
Distillation Network Learning Rate $\{0.00003,0.0003,0.003\}$ $0.0003$
RND Loss Coefficient $\{0.0001,0.001,0.01\}$ $0.01$
RND GAE Coefficient $\{0.01,0.1,1.0\}$ $0.01$
Table 10: Hyperparameters for RND for Craftax-1B and Craftax-1M.
@@ E.2 Craftax-1M
The hyperparameters for Craftax-1M were tuned more thoroughly, as the experiments were very fast to run. We performed a random search over the PPO hyperparameters to obtain a baseline PPO implementation, before then running a random search for ICM and E3B. The hyperparameters for PPO, ICM and E3B are shown in Tables 11, 12 and 13 respectively.